OpenStudy (anonymous):

Translate the problem situation to a system of equations. Do not solve. Bill & Jose's Discount Cinema sold adult tickets for $2.80 each and children's tickets for $1.80 each. Last Tuesday, a total of $71.80 was collected from 36 movie watchers. How many of each type were sold?

OpenStudy (anonymous):

Okay I got this x + y = 4.60; 2.80x + 1.80y = $71.80 if i'm wrong can you explain in steps how you got your answer...

OpenStudy (anonymous):

Good job Esmy! Just be clear when you write the equation for the people. In this case you've assigned children as x, and adults as y. You must write the second equation as: \[x+y=36\]because there were 36 movie watchers (adults and children). Now if you solve this system, you get the following: 7 children and 29 adults.

OpenStudy (anonymous):

Yay so I was right finally -_- lol. Thanks

OpenStudy (anonymous):

Keep up the good work!